Financial data for Joel de Paris, Inc., for last year follow:

Joel de Paris, Inc.
Balance Sheet
Beginning
Balance
Ending
Balance
Assets
Cash $ 125,000 $ 129,000
Accounts receivable 348,000 478,000
Inventory 574,000 488,000
Plant and equipment, net 853,000 845,000
Investment in Buisson, S.A. 393,000 425,000
Land (undeveloped) 251,000 248,000
Total assets $ 2,544,000 $ 2,613,000
Liabilities and Stockholders' Equity
Accounts payable $ 380,000 $ 336,000
Long-term debt 987,000 987,000
Stockholders' equity 1,177,000 1,290,000
Total liabilities and stockholders' equity $ 2,544,000 $ 2,613,000


Joel de Paris, Inc.
Income Statement
Sales $ 4,416,000
Operating expenses 3,621,120
Net operating income 794,880
Interest and taxes:
Interest expense $ 128,000
Tax expense 197,000 325,000
Net income $ 469,880


The company paid dividends of $356,880 last year. The “Investment in Buisson, S.A.,” on the balance sheet represents an investment in the stock of another company. The company's minimum required rate of return of 15%.

Required:

1. Compute the company's average operating assets for last year.

2. Compute the company’s margin, turnover, and return on investment (ROI) for last year. (Round "Margin", "Turnover" and "ROI" to 2 decimal places.)

3. What was the company’s residual income last year?
Answer #1

Average operating assets

= [(2,544,000-251,000-393,000)+(2,613,000-245,000-425,000)]/2

= 1,920,000

Margin = Net operating income/Sales

= 794,880/4,416,000

= 18%

Turnover = Sales/Average operating assets

= 4,416,000/1,920,000

= 2.3

ROI = Margin * Turnover

= 18*2.3

= 41.4%

Residual income = Net operating income - required return

= 794,880 - (1,920,000*15%)

= 506,880